a missile is projected vertically upward from an underground bunker in such a way that t seconds after launch, it's s feet above the ground, where :
s(t)= -16t^2+800t-15.
a) how deep is the bunker?
b) when's the missile at its highest point?
c) what is the max height?
I got the answer of b) is 25s and c) is 9985 but i'm not sure. Thanks if tutors can help me

Answer by nerdybill(7384)

a missile is projected vertically upward from an underground bunker in such a way that t seconds after launch, it's s feet above the ground, where :
s(t)= -16t^2+800t-15.
a) how deep is the bunker?
15 feet
b) when's the missile at its highest point?
at vertex where
t = -b/(2a)
t = -800/(2(-16))
t = -800/(-32)
t = 25 seconds
.
c) what is the max height?
s(t)= -16t^2+800t-15.
s(25)= -16(25)^2+800(25)-15
s(25)= -10000+20000-15
s(25) = 9985 feet